OpenStudy (ray_sihota):

Together Alice, Benita and Carlos collected 150 pounds of Aluminum cans for recycling. Benita collect twice as many pound as carlos while Alice collected 15 POUNDS LESS THEN CARLOS. FIND THE NUMBER OF POUNDS EACH COLLECTED

OpenStudy (campbell_st):

let carlos = x benita = 2x alice = x -15 so x + 2x + x -15 = 150 solve for x, which is the weight carlos collected... then you can find the rest.

OpenStudy (ray_sihota):

X+2X+X-15=150 4X-15=150 4X=165 X=41.25 B=8.250 A=26.25

OpenStudy (campbell_st):

well the equation I wrote was 4x - 15 = 150 solve for x

OpenStudy (ray_sihota):

then add 15 to both sides 4x=165 x=41.25

OpenStudy (campbell_st):

yep... so carlos collected 41.25 lb... based on what's in the question, that's correct

OpenStudy (ray_sihota):

thank you so much the decimal was making me think twice about my answer

OpenStudy (campbell_st):

no because benita is 82.5 and alice is 26.25 they all add to 150